# Enhancing the Accuracy of Spatio-Temporal Models for Wind Speed Prediction by Incorporating Bias-Corrected Crowdsourced Data

### v1

@@ 2.3 Reanalysis Data
In order to better estimate local weather conditions, very high spatial resolution reanalysis products have been created using ERA5 as initial conditions and running weather models at a fine scale. Examples of such datasets include the wind atlas from the Sustainable Energy Authority of Ireland (SEAI, (2013)), and the Global Wind Atlas (GWA) (Davis et al.,, 2023). The SEAI wind atlas is specific to Ireland and designed for wind energy applications, however, it doesn’t contain data at 10m, the operational height of weather stations. In this paper, we focus on the GWA due to its global coverage, and it contains data output at 10m heights. Figure 8 shows the average 10m wind speed modelled by the GWA. The GWA contains static long-term wind speed statistics, such as average wind speed at several heights above ground level, including 10m heights. The spatial resolution of the GWA is a 250m $\times$ 250m grid. Along with average wind speeds on a fine spatial grid, the GWA also provides the shape and scale parameters for the two parameter Weibull distribution on the 250m grid. The Weibull distribution is a frequently used distribution for modelling wind speeds, and will be discussed further in section 3.1. While the spatial resolution is very high, these variables are available only as static, long-term averages; in other words, the GWA does not provide time series or temporal trends that reanalysis datasets such as ERA5 or MERRA2 provide. Previous studies, such as Gruber et al., (2022), have used the GWA to bias correct coarser reanalysis data to create higher resolution maps. In Section 3 we will use GWA reanalysis data to bias correct the crowdsourced stations, while in section 4 the GWA data will provide covariates for our models.
Figure 8: Map of average wind speeds at 10m height, based on data outputted from the Global Wind Atlas. This gives average wind speeds and Weibull parameters on a 250m $\times$ 250m grid (Global Wind Atlas,, 2023).
@@ 3 Bias Correcting Crowdsourced Weather Data
Crowdsourced weather data often exhibits bias compared to measurements from meteorological stations, typically showing a downward bias as discussed in Section 2.2 — a trend also found in previous studies (e.g. Droste et al.,, 2020; Chen et al.,, 2021). After performing the data quality checks described in Section 2.2, we apply a bias correction step before including crowdsourced wind speeds in the spatio-temporal models presented in Section 4.
@@ 3.1 Quantile transformation to the Weibull Distribution
To bias correct the PWS data we utilise a quantile transformation method. Quantile transformations are also commonly used in climate science to bias correct climate models (e.g. Vrac and Vaittinada Ayar,, 2017). Let us consider w, a sample of wind speeds of size $n$ collected at a station, $(w_{i})_{i=1,n}$ with empirical CDF $F_{n}(w)=(\sharp(w_{i})_{i=1,n}\leq w)/n$, for $w\in[0,+\infty)$. The percentile associated to $w_{i}$ is $p_{i}=F_{n}(w_{i})$, and it is given by the rank of $w_{i}$ in the sample divided by $n$. Hence $1/n\leq p_{i}\leq 1$. For our crowdsourced weather data, we calculate empirical CDF of wind speeds separately at each PWS. The percentile value of a datapoint at site $s$ and time $t$ is denoted as $p_{s,t}$. We then assume a theoretical distribution exists for the true wind speed at $s$. As wind speeds are positive, we assume the support of the distribution is $w\in[0,+\infty)$. The CDF of the theoretical true wind speeds is defined as $\hat{F}_{s}\left(w;\Theta\right)$, where $\Theta$ is the vector of parameters characterizing $\hat{F}_{s}$. Given a recorded wind speed $w_{s,t}$ with associated percentile $p_{s,t}$, the theoretical true wind speeds after bias correction can be calculated be performing an inverse CDF transform. Letting $\widetilde{W}_{s,t}$ be the bias corrected wind speeds:
@@
A parametric distribution is required to perform the quantile transformation, as well as an estimate of $\Theta$, the parameter set, at the location of each PWS. Previous studies have examined which probability distribution is best suited to modelling wind speeds (e.g. Carta et al.,, 2009). The most commonly used distribution is the two-parameter Weibull distribution (Wais, (2017) , Bowden et al., (1983)). Therefore, $\hat{f_{s}}\left(w;\Theta\right)$, which bias corrects the PWS is the Weibull distribution. Another reason for choosing the Weibull distribution is that the two parameters are available in the GWA (as discussed in Section 2.3).
@@
[…]
@@
We demonstrate the suitability of a Weibull distribution using four meteorological stations as an illustrative example. Parameters for the Weibull at the four locations are estimated using a maximum likelihood approach (Rockette et al.,, 1974), implemented using the R package EnvStats (Millard,, 2013). Figure 16 in Appendix A shows the QQ plots for these four stations, demonstrating that the Weibull distribution fits well overall, except for some positive outliers. These outliers are due to storm weather and similar extreme events. Future research could consider extensions of that model to extreme values, however this is beyond the scope of this paper.
@@ 3.2 Estimating Weibull Parameters at PWS Locations
The overview of the bias correction approach is included in Figure 9:
@@
The GWA shape ($k$) and scale ($\lambda$) parameters are first downloaded from the online repository (Global Wind Atlas,, 2023). These contain both parameters at 10m on a 250m $\times$ 250m grid. The output is generated using lower resolution reanalysis, and then microscale models are applied using high resolution topographical maps.
@@
To evaluate the accuracy of the parameters from the GWA, we compare them to the Weibull parameters at meteorological stations, which are considered our ground truth. The Weibull parameters at meteorological stations are fit using a maximum likelihood approach, again using the R package EnvStats (Millard,, 2013). In Figure 10 the estimated shape $\{k^{GWA}_{s}\}$ and scale $\{\lambda^{GWA}_{s}\}$ parameters for GWA are compared to the parameters fit at the 23 meteorological stations ($\{k^{MET}_{s},\lambda^{MET}_{s}\}$). Figure 10 includes an identity line for reference. The estimated parameters correlate with the true values but show systematic biases. The GWA underestimates the shape parameter, and overestimates the scale parameter at low-wind sites while remaining accurate at high-wind locations.
@@
Figure 10: Comparison of the Global Wind Atlas (GWA) estimates of the Weibull parameters at the 23 Met Éireann weather station locations. (a) The GWA systematically underestimates the Weibull shape parameter. (b) The GWA overestimates the scale parameter at low-wind sites but is accurate for high-wind locations. The blue line in each plot represents the identity line, $y=x$.
To investigate if the accuracy of the scale parameter is spatially dependent, a map of the errors for the GWA parameters is shown in Figure 11. Studying the GWA plot we can see the model has good accuracy is coastal areas, however it significantly overestimates wind speed in inland locations.
Figure 11: The errors of GWA estimate of Weibull scale parameter, highlighting increased errors in inland locations.
We consider calibration models, $h^{k}(\cdot)$ and $h^{\lambda}(\cdot)$, which describe the relationship between the GWA shape and scale parameters at each location and the true values derived from meteorological station observations. Our baseline models were two simple linear regressions with $\{k^{GWA}_{s},\lambda^{GWA}_{s}\}$ treated as covariates, and $\{k^{MET}_{s},\lambda^{MET}_{s}\}$ as dependent variables. The accuracy of each calibration model was compared using a leave-one out framework, to test how models generalise to unobserved stations. For our shape parameter a linear regression was used as $h^{k}(\cdot)$, the final calibration model: $k^{MET}_{s}=\beta_{0}+\beta_{1}k^{GWA}_{s}+\epsilon^{k}_{s}$, where $\epsilon^{k}_{s}\sim\mathcal{N}(0,\sigma^{2}_{k})$ is independent Gaussian noise.
For the scale parameter, two extensions beyond linear regression are included. To model non-linearity in the relationship between $\lambda^{GWA}_{s}$ and $\lambda^{MET}_{s}$, a generalized additive model (GAM) was implemented which used a smooth spline term. Thin plate splines provide a flexible way of modelling non-linear relationships, while also controlling smoothness to prevent overfitting. To incorporate the dependence seen in Figure 11 distance from the see is added as a linear covariate. The final calibration model, $h^{\lambda}(\cdot)$, for the scale parameter is given as: $h^{\lambda}(\cdot):\quad\hat{\lambda}^{PWS}_{s}=\beta_{0}+s\left(\lambda^{GWA}_{s}\right)+\beta_{1}\left(\text{distance from sea (km)}\right)+\epsilon^{\lambda}_{s}$, where $\epsilon^{\lambda}_{s}\sim\mathcal{N}(0,\sigma^{2}_{\lambda})$ is independent Gaussian noise.
The full results for the different models are included in Appendix B, along with estimated coefficients.
@@
For the shape parameter, a simple linear regression relates the GWA estimate at each location to the predicted value at the PWS.
[…]
For the scale parameter, $s(\lambda^{GWA}_{s})$ is a thin plate spline function of the GWA scale estimate, while the distance from the sea is included as a linear term. These models produce a predicted Weibull distribution for wind speed at each PWS.
@@
Since there is no ground truth at PWS, the accuracy of bias correction cannot be quantified directly. However, Section 5 will demonstrate the benefits of incorporating bias-corrected stations into spatial models. To visualize the impact of the corrections, Figure 12 examines the four station pairs from Figure 5, where the corrected wind speeds show closer agreement with neighbouring meteorological stations.
@@
To compare the distribution of corrected wind speeds and raw wind speeds at a sample site, the PWS labelled 1 in Figure 4 is chosen, and a kernel density estimate of raw wind speeds and corrected wind speeds are shown in Figure 13. The corrected distribution shifts the wind speed distribution to the right. As expected, the quantile transformation results in wind speeds which closely follows a Weibull distribution. Note that while the raw wind speeds appear to have a slight bimodal distribution, this could be due to PWS often storing only discrete values or exhibiting zero inflation. One potential limitation of the quantile transformation is its tendency to underestimate very high wind speeds, particularly those with negligible probabilities under the Weibull distribution. However, in the context of wind energy, this is less critical since turbines shut down at extreme wind speeds (Hering and Genton,, 2010). Based on visual verification, the quantile transformation generally produces wind speeds at PWS that closely match those from neighbouring meteorological stations.

### v2

@@ 2.3 Reanalysis Data
In order to better estimate local weather conditions, very high spatial resolution reanalysis products have been created using ERA5 as initial conditions and running weather models at a fine scale. Examples of such datasets include the wind atlas from the Sustainable Energy Authority of Ireland (SEAI,, 2013), and the Global Wind Atlas (GWA) (Davis et al.,, 2023). The SEAI wind atlas is specific to Ireland and designed for wind energy applications, however, it doesn’t contain data at 10m, the operational height of weather stations. In this paper, we focus on the GWA due to its global coverage, and it contains data output at 10m heights. Figure 8 shows the average 10m wind speed modelled by the GWA. The GWA contains static long-term wind speed statistics, such as average wind speed at several heights above ground level, including 10m heights. The spatial resolution of the GWA is a 250m $\times$ 250m grid. Along with average wind speeds on a fine spatial grid, the GWA also provides the shape and scale parameters for the two parameter Weibull distribution on the 250m grid. The Weibull distribution is a frequently used distribution for modelling wind speeds, and will be discussed further in section 3.2. While the spatial resolution is very high, these variables are available only as static, long-term averages; in other words, the GWA does not provide wind speed time series as is the case for the ERA5 and MERRA2 reanalysis datasets, but instead GWA only provides statistics that characterize the probability distribution of wind speed at each grid point and target height. Previous studies, such as Gruber et al., (2022), have used the GWA to bias correct coarser reanalysis data to create higher resolution maps. In Section 3 we will use GWA reanalysis data to bias correct the crowdsourced stations, while in section 4 the GWA data will provide covariates for our models.
Figure 8: Map of average wind speeds at 10m height on a 250m $\times$ 250m grid, from the Global Wind Atlas (Global Wind Atlas,, 2023).
@@ 3 Bias Correcting Crowdsourced Weather Data
Crowdsourced weather data often exhibits bias compared to measurements from meteorological stations, typically showing a downward bias as discussed in Section 2.2 — a trend also found in previous studies (e.g. Droste et al.,, 2020; Chen et al., 2021c, ). After performing the data quality checks described in Section 2.2, we apply a bias correction step before including crowdsourced wind speeds in the spatio-temporal models presented in Section 4.
@@
3.1 Selecting optimal distribution
To perform the quantile transformation, we examine several parametric distributions to identify the most appropriate for modelling Irish wind speeds. Commonly used distributions in wind speed modelling include the Weibull, log-normal, Gamma, and Rayleigh (e.g. Jung and Schindler,, 2019; Chen et al., 2021a, ; Ouarda et al.,, 2015). We focus on the first three distributions, as the Rayleigh is a special case of the Weibull with shape parameter $k=2$. Table 2 provides the probability density functions and parameters for each distribution.
Distribution Probability Density Function Parameters
Weibull $f(x)=\frac{k}{\lambda}\left(\frac{x}{\lambda}\right)^{k-1}e^{-(x/\lambda)^{k}}$ $k>0$ (shape), $\lambda>0$ (scale)
Gamma $f(x)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-\beta x}$ $\alpha>0$ (shape), $\beta>0$ (rate)
Log-normal $f(x)=\frac{1}{x\sigma\sqrt{2\pi}}\exp\left(-\frac{(\ln x-\mu)^{2}}{2\sigma^{2}}\right)$ $\mu\in\mathbb{R}$ (log-mean), $\sigma>0$ (log-sd)
Table 2: Probability distributions considered for wind speed modelling
Following previous studies (Jung and Schindler,, 2019; Chen et al., 2021a, ; Ouarda et al.,, 2015), we first compare maximised log-likelihoods for each fitted distribution at each station. As each distribution contains two parameters, a penalised criteria such as AIC is not required. The number of stations where each distribution has the highest log-likelihood is recorded. Second, we use the Kolmogorov-Smirnov test (Berger and Zhou,, 2014) to measure maximum distance between empirical and theoretical cumulative distribution functions. Finally, we assess tail behaviour by evaluating the absolute difference between the empirical and theoretical 95th percentiles. Parameters for each distribution are estimated using a maximum likelihood approach (Rockette et al.,, 1974), implemented using the R package fitdistrplus (Delignette-Muller and Dutang,, 2015). Figure 9 illustrates the three fitted distributions at one representative site.
Figure 9: Three potential distributions along with histogram of wind speeds at Mace head (Site 1 in Figure 4).
Test Weibull Gamma Log-normal
Highest log-likelihood (count) 11 12 0
Kolmogorov-Smirnov distance (mean) 0.031 0.044 0.076
95th percentile absolute difference (m/s) (mean) 0.111 0.418 1.412
Table 3: Goodness of fit for candidate distributions (mean values across 23 stations). For log-likelihood, higher values indicate better performance; for distance and difference metrics, lower values indicate better performance.
Table 3 shows performance for the three distributions, summarised across 23 meteorological stations. From this we see that, for the log-likelihood, the Weibull and Gamma distributions are highest at a similar number of stations. Furthermore, the Weibull distribution outperforms the other two distributions for average KS statistic and also provides the most accurate representation of the 95th percentile. Overall this demonstrates superior overall fit and tail behaviour for the Weibull.
In addition, the Weibull distribution is widely used in wind resource assessment and is directly available in reanalysis products such as the Global Wind Atlas, making it practical for bias correction. This is consistent with previous studies advocating the Weibull distribution for wind speed modelling (Carta et al.,, 2009; Wais,, 2017; Bowden et al.,, 1983; Ouarda et al.,, 2015).
To further illustrate the suitability of the Weibull distribution, Appendix A presents QQ-plots for four representative meteorological stations, showing generally good agreement with the empirical distribution except for several positive outliers.
@@ 3.2 Quantile transformation to the Weibull distribution
To bias correct the PWS data we utilise a quantile transformation method. Quantile transformations are also commonly used in climate science to bias correct climate models (e.g. Vrac and Vaittinada Ayar,, 2017). Let us consider w, a sample of wind speeds of size $n$ collected at a station, $(w_{i})_{i=1,n}$ with empirical CDF $F_{n}(w)=(\sharp(w_{i})_{i=1,n}\leq w)/n$, for $w\in[0,+\infty)$. The percentile associated to $w_{i}$ is $p_{i}=F_{n}(w_{i})$, and it is given by the rank of $w_{i}$ in the sample divided by $n$. Hence $1/n\leq p_{i}\leq 1$. For our crowdsourced weather data, we calculate empirical CDF of wind speeds separately at each PWS. The percentile value of a datapoint at site $s$ and time $t$ is denoted as $p_{s,t}$. We then assume a theoretical distribution exists for the true wind speed at $s$. As wind speeds are positive, we assume the support of the distribution is $w\in[0,+\infty)$. The CDF of the theoretical true wind speeds is defined as $\hat{F}_{s}\left(w;\Theta\right)$, where $\Theta$ is the vector of parameters characterizing $\hat{F}_{s}$. Given a recorded wind speed $w_{s,t}$ with associated percentile $p_{s,t}$, the bias-corrected wind speeds can be calculated by performing an inverse CDF transform. Letting $\widetilde{W}_{s,t}$ be the bias corrected wind speeds:
@@
A parametric distribution is required to perform the quantile transformation, as well as an estimate of $\Theta$, the parameter set, at the location of each PWS. As found in Section 3.1, the Weibull distribution most closely models the measured wind speeds at Irish meteorological stations. Therefore, $\hat{f_{s}}\left(w;\Theta\right)$, which bias corrects the PWS is the Weibull distribution.
@@
where $\Gamma\left(\cdot\right)$ denotes the Gamma function, $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\,dt$. In order to carry out the quantile transformation bias correction on wind speeds, the parameters $\lambda$ and $k$ of the Weibull PDF need to be estimated at the location of each PWS. The estimated parameters are used to transform empirical percentiles to corrected wind speeds using an inverse CDF transform. For a Weibull distribution, the inverse CDF of a probability $p\in[0,1)$ defines a quantile given by inverting Equation (3):
@@ 3.3 Estimating Weibull Parameters at PWS Locations
Selecting a Suitable Reanalysis Dataset
To bias correct the PWS, we require Weibull shape ($k$) and scale ($\lambda$) parameters at each station location. Reanalysis products provide complete spatial coverage over the study area, enabling parameter estimation at PWS sites. We compare two reanalysis datasets: ERA5, which provides historical wind speed time series and was previously shown to be the most accurate for Ireland (Doddy Clarke et al.,, 2021), and the Global Wind Atlas (GWA), which offers higher spatial resolution (250m) and directly outputs Weibull parameters.
For each dataset, Weibull parameters are obtained at all meteorological station locations. For the GWA, published Weibull parameters are directly downloaded. These are interpolated to the station locations using inverse-distance weighting. For ERA5, one year of hourly wind speeds preceding the study period is downloaded. Weibull parameters are estimated at each ERA5 grid cell using maximum likelihood and then interpolated to the station locations. Table 4 reports the correlation between the reanalysis-derived parameters and the station-derived parameters.
Metric GWA vs. Stations ERA5 vs. Stations
Shape parameter correlation ($k$) 0.46 0.23
Scale parameter correlation ($\lambda$) 0.91 0.78
Table 4: Correlation between reanalysis Weibull parameters and those derived from meteorological station observations.
The GWA has substantially stronger agreement with meteorological stations for the scale parameter and moderately better correlation for the shape parameter. Although both datasets show limited skill for estimating the shape parameter, due to lower spatial variability, the GWA’s higher resolution provides a closer match overall. For this reason, the GWA is selected as the distribution for bias correction. Although high correlation does not imply that the GWA parameters match the station‐derived values in absolute magnitude, it indicates a strong linear association. This, in turn, supports the feasibility of learning a calibration function that maps GWA Weibull parameters to those estimated at meteorological stations.
The full overview of the bias correction approach is displayed in Figure 10:
@@
The GWA shape ($k$) and scale ($\lambda$) parameters are first downloaded from the online repository (Global Wind Atlas,, 2023). These contain both parameters at 10m height on a 250m $\times$ 250m grid.
@@
To evaluate the accuracy of the parameters from the GWA, we compare them to the Weibull parameters estimated from wind speed time series recorded at meteorological stations, which are considered our ground truth. The Weibull parameters at meteorological stations are estimated using a maximum likelihood approach using the R package fitdistrplus (Delignette-Muller and Dutang,, 2015). Figure 11 compares the shape $\{k^{GWA}_{s}\}$ and scale $\{\lambda^{GWA}_{s}\}$ parameters from GWA to the parameters estimated at the 23 meteorological stations ($\{k^{MET}_{s},\lambda^{MET}_{s}\}$). Results show that the parameters derived from GWA correlate well with the true values but show systematic biases. The GWA underestimates the shape parameter, and overestimates the scale parameter at low-wind sites while remaining accurate at high-wind locations.
@@
Figure 11: Comparison of the parameters of the Weibull distribution derived from the GWA and from observations at the 23 Met Éireann weather stations. (a) Weibull shape parameter. (b) Weibull scale parameter. The blue line in each plot represents the identity line, $y=x$.
To investigate whether the accuracy of the Weibull parameters is spatially dependent, Figure 12 maps the residuals for both the shape and scale parameters. The shape parameter shows no clear spatial structure in its errors. In contrast, the scale parameter exhibits a mild spatial pattern: coastal locations generally display low error, whereas larger errors tend to occur at inland stations. While this pattern is not uniform across all sites, it motivates the inclusion of distance from the coastline as an explanatory variable in the calibration model.
(a) Shape parameter residuals
(b) Scale parameter residuals
Figure 12: Spatial distribution of the errors (Truth - Prediction) in GWA-derived Weibull parameters. (a) Shape parameter residuals; (b) Scale parameter residuals.
To mitigate the errors of GWA estimates of the Weibull scale and shape parameters, we consider two calibration models, $h^{k}(\cdot)$ and $h^{\lambda}(\cdot)$, which describe the relationship between the GWA parameters at each location and the true values derived from meteorological station observations.
According to the structure of errors observed in Figure 11(a), and the lack of spatial patterns in Figure 12(a), we selected a simple regression model for the shape parameter, in which $k^{GWA}_{s}$ is treated as a covariate and $k^{MET}_{s}$ as the dependent variable. The calibration model for the shape parameter is therefore: $h^{k}(\cdot):\hat{k}^{MET}_{s}=\beta_{0}+\beta_{1}k^{GWA}_{s}+\epsilon^{k}_{s}$, where $\epsilon^{k}_{s}\sim\mathcal{N}(0,\sigma^{2}_{k})$ is independent Gaussian noise.
For the scale parameter, the structure of errors observed in Figure 11(b) and its spatial pattern evidenced in Figure 12(b) made us select a generalized additive model (GAM) with two covariates: $k^{GWA}_{s}$ (with thin plate splines as link function) and the distance from the sea (with a linear link). The calibration model for the scale parameter is therefore: $h^{\lambda}(\cdot):\quad\hat{\lambda}^{MET}_{s}=\beta_{0}+s\left(\lambda^{GWA}_{s}\right)+\beta_{1}\left(\text{distance from sea (km)}\right)+\epsilon^{\lambda}_{s}$, where $\epsilon^{\lambda}_{s}\sim\mathcal{N}(0,\sigma^{2}_{\lambda})$ is independent Gaussian noise.
More details on calibration models selection and estimated coefficients are included in Appendix B. To assess the suitability of the calibrated Weibull distribution derived from the GWA for bias correction, we compared the implied distributional properties against empirical wind speed statistics at the meteorological stations. Three aspects of the wind speed distribution were evaluated: (i) the mean, (ii) the variance, and (iii) tail behaviour, assessed using the 95th percentile.
Metric Mean Absolute Error (MAE) Pearson Correlation
Mean Wind Speed 0.38 0.94
Variance 0.61 0.95
95th Percentile 0.65 0.94
Table 5: Comparison between empirical wind speed statistics and those implied by the calibrated Weibull distribution across stations. Metrics are aggregated over the 23 meteorological stations.
Table 5 highlights that the calibrated Weibull distributions accurately capture key characteristics of the observed wind speed distributions. Across stations, the implied means, variances, and 95th percentiles show high correlation (0.94–0.95) with empirical values and low mean absolute errors. This indicates that the Weibull representation is consistent with the observed wind speeds and suitable as a target distribution for the bias-correction procedure.
@@
$h^{\lambda}(\cdot):\quad\hat{\lambda}^{PWS}_{s}=\beta_{0}+s\left(\lambda^{GWA}_{s}\right)+\beta_{1}\left(\text{Distance from sea (km)}\right)$ (7)
@@
Since there is no ground truth at PWS, the accuracy of bias correction cannot be quantified directly. However, Section 5 will demonstrate the benefits of incorporating bias-corrected stations into spatial models. To visualize the impact of the corrections, Figure 13 examines the four pairs of stations from Figure 5, where the corrected wind speeds show closer agreement with neighbouring meteorological stations.
@@
To compare the distribution of bias-corrected and raw wind speeds at a sample site, the PWS labelled 1 in Figure 4 is chosen, and a kernel density estimate of raw wind speeds and corrected wind speeds are shown in Figure 14. The corrected distribution shifts the wind speed distribution to the right. As expected, the quantile transformation results in wind speeds which closely follows a Weibull distribution. As the quantile transformation maps the empirical percentiles to a parametric distribution, isolated extreme values in the raw data have reduced impact on the overall model. Outlying raw observations are still mapped to high percentiles, but their absolute magnitude becomes consistent with the overall fitted distribution.
Note that while the raw wind speeds appear to have a slight bimodal distribution, this could be due to PWS often storing only discrete values or exhibiting zero inflation. One potential limitation of the quantile transformation is its tendency to underestimate very high wind speeds, particularly those with negligible probabilities under the Weibull distribution. The impact of extremes is further discussed in the Section 5.2. Based on visual verification, the quantile transformation generally produces wind speeds at PWS that closely match those from neighbouring meteorological stations.